OpenStudy (anonymous):

The length of a rectangle is 5/6 feet. The width is 3/8 feet, How much greater is the length of the rectangle than the width?

OpenStudy (anonymous):

(5/6)-(3/8) = (8*5-6*3)/(6*8) = 22/48 = 11/24 ft.

OpenStudy (anonymous):

metal

OpenStudy (anonymous):

\[\frac{ 5 }{ 6}-\frac{ 3 }{ 8 },L.C.M.~ of~ 6~ and~ 8~ is~ 24\] \[\frac{ 5*4-3*3 }{ 24 }=\frac{ 20-9 }{24 }=\frac{ 11 }{ 24 }\]